# Demystifying amortized causal discovery with transformers

### v1

@@ A.2 CSIvA architecture
Transformers [25] are a popular neural network architecture for modeling structured, sequential data data. They consist of an encoder, a stack of layers that learns a representation of each element in the input sequence based on its relation with all the other sequence’s elements, through the mechanism of self-attention, and a decoder, which maps the learned representation to the target of interest. Note that data for causal discovery are not sequential in their nature, which motivates the adaptations introduced by Ke et al. [1] in their CSIvA architecture.
@@
Similarly to the transformer’s encoder, CSIvA stacks a number of identical layers, performing self-attention followed by a nonlinear mapping, most commonly an MLP layer. The main difference relative to the standard encoder is in the implementation of the self-attention layer: as transformers are in their nature suitable for the representation of sequences, given an input sample of $D$ elements, self-attention is usually run across all elements of the sequence. However, data for causal discovery are tabular, rather than sequential: one option would be to unravel the $n\times d$ matrix of the data, where $n$ is the number of observations and $d$ the number of variables, into a vector of $n\cdot d$ elements, and let this be the input sequence of the encoder. CSIvA adopts a different strategy: the self-attention in each encoder layer consists of alternate passes over the attribute and the sample dimensions, known as alternating attention [37]. As a clarifying example, consider a dataset $\{(x_{1}^{i},x_{2}^{i})\}_{i=1}^{n}$ of $n$ i.i.d. samples from the joint distribution of the pair of random variables $X_{1},X_{2}$. For each layer of the encoder, in the first step (known as attention between attributes), attention operates across all nodes of a single sample $(x_{1}^{i},x_{2}^{i})$ to encode the relationships between the two nodes. In the second step (attention between samples), attention operates across all samples $(x_{k}^{1},\ldots,x_{k}^{n}),k\in\{1,2\}$ of a given node, to encode information about the distribution of single node values.
@@
The decoder uses the summary information from the encoder to generate a prediction of the adjacency matrix $A$ of the underlying $\mathcal{G}$. It operates sequentially, at each step producing a binary output indicating the prediction $\hat{A}_{i,j}$ of $A_{i,j}$, proceeding row by row. The decoder is an autoregressive transformer, meaning that each prediction $\hat{A}_{i,j}$ is obtained based on all elements of $A$ previously predicted, as well as the summary produced by the encoder. The method does not enforce acyclicity, although Ke et al. [1] shows that in cyclic outputs genereally don’t occur, in practice.
@@ B.2 Synthetic data
In this section, we provide additional details on the synthetic data generation, which was performed with the causally [^2] https://causally.readthedocs.io/en/latest/ Python library [9]. Our data-generating framework follows that of Montagna et al. [9], an extensive benchmark of causal discovery methods on different classes of SCMs.
Causal mechanisms.
The nonlinear mechanisms of the PNL model and the nonlinear ANM model are generated by a neural network with one hidden layer with $10$ hidden units, with a parametric ReLU activation function. The network weights are randomly sampled according to a standard Gaussian distribution. The linear mechanisms are generated by sampling the regression coefficients in the range $[-3,-0.5]\cup[0.5,3]$.
@@
Data are standardized with their empirical variance, which removes the presence of shortcuts which could be learned by the network, notably varsortability [32] and score-sortability [33].
@@ B.3 Computer resources
Appendix C Further experiments
@@
We present our experimental results on one further question, to help clarify the results in the main text of the paper. Our aim is to understand when to make tradeoffs between computational resources, and having models that have been trained on a wider variety of SCMs. We compare training on multiple SCMs to single-SCM training, when all models see the same amount of training data from each SCM type as a non-mixed model (i.e. a mixed network trains on $15,000$ linear datasets and $15,000$ PNL datasets, instead of $15,000$ divided between the two SCM types).
@@
In the main text of this paper, we compare neural networks trained on a mix of structural causal models (e.g. noise distributions, or mechanism types), to models trained on a single mechanism-noise combination, where all models have the same amount of training data, $15,000$ datasets. In mixed training, we split these evenly, so a "lin, nl" model is trained on $7,500$ datasets from linear SCMs, and $7,500$ from nonlinear SCMs. Our results in this framework are promising, and show that for many combinations of SCM types, we can train one model instead of two, and achieve good progress, while making a $50\%$ savings on training costs. However, if our training budget is high/unlimited, we should also ask whether there is a downside to mixed training - can we achieve the same performance as a model trained on a single SCM type? Fig. 6 shows good results in this direction - the models trained with the same number of datasets per SCM type as an unmixed model had similar (or even better, for PNL data) performance as the un-mixed model trained on the same SCM type as the test data. These mixed models are also significantly more useful than having 2 or 3 separate models per SCM type, as they have good across-the-board performance. However, if we used the same computational resources to train 3 separate networks (one for each mechanism type) and wanted to use them for causal discovery on a dataset with unknown assumptions, we would be left with the rather difficult task of deciding which model to trust.
@@
Appendix D Theoretical results and proofs

### v2

@@ A.2 CSIvA architecture
Transformers (Vaswani et al., 2017) are a popular neural network architecture for modeling structured, sequential data data. They consist of an encoder, a stack of layers that learns a representation of each element in the input sequence based on its relation with all the other sequence’s elements, through the mechanism of self-attention, and a decoder, which maps the learned representation to the target of interest. Note that data for causal discovery are not sequential in their nature, which motivates the adaptations introduced by Ke et al. (2023b) in their CSIvA architecture.
@@
Similarly to the transformer’s encoder, CSIvA stacks a number of identical layers, performing self-attention followed by a nonlinear mapping, most commonly an MLP layer. The main difference relative to the standard encoder is in the implementation of the self-attention layer: as transformers are in their nature suitable for the representation of sequences, given an input sample of $D$ elements, self-attention is usually run across all elements of the sequence. However, data for causal discovery are tabular, rather than sequential: one option would be to unravel the $n\times d$ matrix of the data, where $n$ is the number of observations and $d$ the number of variables, into a vector of $n\cdot d$ elements, and let this be the input sequence of the encoder. CSIvA adopts a different strategy: the self-attention in each encoder layer consists of alternate passes over the attribute and the sample dimensions, known as alternating attention Kossen et al. (2021). As a clarifying example, consider a dataset $\{(x_{1}^{i},x_{2}^{i})\}_{i=1}^{n}$ of $n$ i.i.d. samples from the joint distribution of the pair of random variables $X_{1},X_{2}$. For each layer of the encoder, in the first step (known as attention between attributes), attention operates across all nodes of a single sample $(x_{1}^{i},x_{2}^{i})$ to encode the relationships between the two nodes. In the second step (attention between samples), attention operates across all samples $(x_{k}^{1},\ldots,x_{k}^{n}),k\in\{1,2\}$ of a given node, to encode information about the distribution of single node values.
@@
The decoder uses the summary information from the encoder to generate a prediction of the adjacency matrix $A$ of the underlying $\mathcal{G}$. It operates sequentially, at each step producing a binary output indicating the prediction $\hat{A}_{i,j}$ of $A_{i,j}$, proceeding row by row. The decoder is an autoregressive transformer, meaning that each prediction $\hat{A}_{i,j}$ is obtained based on all elements of $A$ previously predicted, as well as the summary produced by the encoder. The method does not enforce acyclicity, although Ke et al. (2023b) shows that in cyclic outputs genereally don’t occur, in practice.
@@ B.2 Synthetic data
In this section, we provide additional details on the synthetic data generation, which was performed with the causally [^2] https://causally.readthedocs.io/en/latest/ Python library (Montagna et al., 2023a). Our data-generating framework follows that of Montagna et al. (2023a), an extensive benchmark of causal discovery methods on different classes of SCMs.
@@
Causal mechanisms.
The nonlinear mechanisms of the PNL model and the nonlinear ANM model are generated by a neural network with one hidden layer with $10$ hidden units, with a parametric ReLU activation function. The network weights are randomly sampled according to a standard Gaussian distribution (we refer to data with nonlinear mechanisms sampled according to this approach as NN-data). The linear mechanisms are generated by sampling the regression coefficients in the range $[-3,-0.5]\cup[0.5,3]$.
NN-data generation: literature review.
We present an extensive list of works adopting neural networks for the sampling of nonlinear mechanisms, similarly to our work: Brouillard et al. (2020; 2021), Lippe et al. (2022), Bello et al. (2022), Montagna et al. (2023a; b), Ke et al. (2023a; b), Reizinger et al. (2023), Massidda et al. (2023), Tran et al. (2024). This suggests that our data generation strategy is established in the literature of causality. Additional experiments with sampling of nonlinear mechanisms from Gaussian processes are presented in Appendix D.7.
Data are standardized with their empirical variance, which removes the presence of shortcuts which could be learned by the network, notably varsortability (Reisach et al., 2021) and score-sortability (Montagna et al., 2023b).
@@ B.3 Computer resources
Appendix C CSIvA identifiability properties on multivariate SCMs
In the main manuscript, we limit our empirical and theoretical analysis of the identifiability guarantees provided by CSIvA to the case of bivariate causal models. In this section, we show how our findings are expected to extend to the multivariate setting. Our starting point is Theorem 28 from Peters et al. (2014): the intuition is that identifiability of multivariate additive noise models can be guaranteed by iteratively verifying that the causal order of all bivariate subgraphs is individually identifiable. We formalize this reporting the following set of definitions and results from Peters et al. (2014).
Condition 1 (Condition 19 of Peters et al. (2014)).
Consider an additive noise model with structural equations $X_{2}=f(X_{1})+N$, $X_{1},N$ independent random variables. The triple $(f,p_{X_{1}},p_{N})$ does not solve the following differential equation for all pairs $x_{1},x_{2}$ with $f^{\prime}(x_{2})\nu^{\prime\prime}(x_{2}-f(x_{1}))\neq 0$:
$\xi^{\prime\prime\prime}=\xi^{\prime\prime}\left(\frac{f^{\prime\prime}}{f^{\prime}}-\frac{\nu^{\prime\prime\prime}f^{\prime}}{\nu^{\prime\prime}}\right)+\frac{\nu^{\prime\prime\prime}\nu^{\prime}f^{\prime\prime}f^{\prime}}{\nu^{\prime\prime}}-\frac{\nu^{\prime}(f^{\prime\prime})^{2}}{f^{\prime}}-2\nu^{\prime\prime}f^{\prime\prime}f^{\prime}+\nu^{\prime}f^{\prime\prime\prime},$ (9)
Here, $\xi\coloneqq\log p_{X_{1}}$, $\nu\coloneqq\log p_{N}$, the logarithms of the strictly positive densities. The arguments $x_{2}-f(x_{1})$, $x_{1}$, and $x_{1}$ of $\nu$, $\xi$ and $f$ respectively, have been removed to improve readability.
The intuition is that a bivariate additive noise model (which can be seen as a reparametrization of a post-nonlinear model, as shown in our Section 4.4) is identifiable if it has a density that satisfies the above Condition 1. This can be generalized to the case of multivariate ANMs where, for identifiability to hold, Condition 1 must be verified for each pair of causally related variables in the SCM: when this is verified, we refer to a restricted additive noise model.
Definition 2 (Definition 27 of Peters et al. (2014)).
Consider an additive noise model with structural equations $X_{i}\coloneqq f_{i}(X_{\operatorname{PA}^{\mathcal{G}}_{i}})+N_{i},i=1,\ldots,k$, independent noise terms and causal graph $\mathcal{G}$. We call this SCM a restricted additive noise model if for all $X_{j}\in X$, $X_{i}\in X_{\operatorname{PA}^{\mathcal{G}}_{j}}$, and all sets $X_{S}\subseteq X$, $S\subset\mathbb{N}$, with $X_{\operatorname{PA}^{\mathcal{G}}_{j}}\setminus\{X_{i}\}\subseteq X_{S}\subseteq X_{\operatorname{ND}_{j}}^{\mathcal{G}}\setminus\{X_{i},X_{j}\}$ (where $\operatorname{ND}_{j}$ is the set of non descendants of the node $X_{j}$ in the graph $\mathcal{G}$), there is a value $x_{S}$ with $p(x_{S})>0$, such that the triplet
$(f_{j}(x_{\operatorname{PA}^{\mathcal{G}}_{j}\setminus\{i\}},\cdot),p_{X_{i}|X_{S}=x_{S}},p_{N_{j}})$
satisfies Condition 1. Here, $f_{j}(x_{\operatorname{PA}^{\mathcal{G}}_{j}\setminus\{i\}},\cdot)$ denotes the mechanism function $x_{i}\mapsto f_{j}(x_{\operatorname{PA}^{\mathcal{G}}_{j}})$. Additionally, we require the noise variables to have positive densities and the functions $f_{j}$ to be continuous and three times continuously differentiable.
In the above definition we adopted the following notation: for a random vector $X=(X_{1},\ldots,X_{n})$, and a set $S\subseteq\{1,\ldots,n\}$, we define $X_{S}$ as the vector with elements $\{X_{i}:i\in S\}$.
Finally, the next theorem formalizes the intuition we’ve advocated so far: the restricted additive noise model of Definition 2, i.e. an SCM whose pairwise causal relations are individually identifiable, is itself identifiable.
Theorem 1 (Theorem 28 of Peters et al. (2014)).
Let $X$ be generated by a restricted additive noise model with graph $\mathcal{G}$, and assume that the causal mechanisms $f_{j}$ are not constant in any of the input arguments, i.e. for $X_{i}\in X_{\operatorname{PA}^{\mathcal{G}}_{j}}$, there exist $x_{i}\neq x^{\prime}_{i}$ such that $f_{j}(x_{\operatorname{PA}^{\mathcal{G}}_{j}\setminus\{i\}},x_{i})\neq f_{j}(x_{\operatorname{PA}^{\mathcal{G}}_{j}\setminus\{i\}},x^{\prime}_{i})$. Then, $\mathcal{G}$ is identifiable.
We note that this relation between bivariate and multivariate identifiability was recently exploited for causal discovery with optimal transport by Tu et al. (2022)
Discussion and multivariate identifiability guarantees of transformers.
The above theorem states that a restricted ANM is identifiable. According to our Definition 2, an additive noise model $X_{h}\coloneqq f_{h}(X_{\operatorname{PA}^{\mathcal{G}}_{h}})+N_{h},h=1,\ldots,k$ is restricted if each pair of connected nodes $X_{i}\rightarrow X_{j}$, for each $X_{\operatorname{PA}^{\mathcal{G}}_{j}}\setminus\{X_{i}\}\subseteq X_{S}\subseteq X_{\operatorname{ND}_{j}}^{\mathcal{G}}\setminus\{X_{i},X_{j}\}$ (think of $X_{S}$ as the set of all possible causes of $X_{j}$, except $X_{i}$), can define a bivariate SCM of the form $(f_{j}(x_{\operatorname{PA}^{\mathcal{G}}_{j}\setminus\{i\}},\cdot),p_{X_{i}|X_{S}=x_{S}},p_{N_{j}})$ that satisfies Condition 1, i.e. that is identifiable. How does this relate to our findings? Our experiments and analysis of Section 4.3 validate the hypothesis that transformers align with the theory of identifiability in the case of training and inference on bivariate graphs. Given Theorem 1, we know that multivariate identifiability is a property of SCMs where each pair of causes and effects can define a bivariate structural causal model that is itself identifiable: this implies that the empirical guarantees of identifiability we verify for transformers (via CSIvA) on bivariate models must extend to multivariate models. This is apparent by contradiction: say we train a CSIvA architecture that can infer the causal direction of a multivariate linear Gaussian model (which is notoriously non-identifiable). This means that our algorithm can infer the causal direction for each bivariate subgraph consisting of two variables connected according to a linear Gaussian structural equation: this would contradict our experimental results presented in Fig. 3(b) and analyzed in Section 4.3.
Appendix D Further experiments
In this section, we provide additional experiments on real-world data, on the scaling properties of CSIvA in the number of training samples, and benchmark CSIvA performance in comparison to several well-established or state-of-the-art methods for causal discovery, with identifiability guarantees under different assumptions: DirectLiNGAM (Shimizu et al., 2011) for inference on linear non-Gaussian models, CAM (Bühlmann et al., 2014) for inference of additive noise models with additive mechanisms, NoGAM (Montagna et al., 2023c) and GraNDAG (Lachapelle et al., 2020) for inference on ANMs (the latter is taken from the continuous optimization literature of causal discovery, already mentioned in the related works Section 2).
D.1 Experiments on real-world datasets
We consider the accuracy of CSIvA trained on different dataset configurations and tested on real-world datasets. In particular, we perform evaluation on the Tübingen pairs dataset (Mooij et al., 2016), the Sachs biological dataset (Sachs et al., 2005), the AutoMPG dataset on cars fuel consumption (Bache & Lichman, 2013) and the Sprinkler dataset, a simple dataset on the causal relations between the binary categorical variables rain, sprinkler on/off, wet grass. Given that our algorithms are trained on bivariate models, from each multivariate dataset we extract all possible two variables subgraphs where this operation does not introduce new confounding effects. This results in $9$ datasets from Sachs, $3$ datasets from AutoMPG, and $2$ datasets from Sprinkler. We consider $102$ pairs from the Tübingen dataset.
Real-world generalization of mixed-trained models.
In Fig. 6 we illustrate the average accuracy per dataset type (Sachs, AutoMPG, Sprinkler, Tübingen) of each CSIvA model. In particular, we want to probe the goodness of mixed training in real-world scenarios. To this end, we train four architectures on the following dataset configurations: linear-mlp, anm-mlp, pnl-mlp, mixed-mixed, where the latter denotes the model trained on SCMs with linear, additive nonlinear and post-nonlinear mechanisms, and Beta, Gamma, Gumbel, Exponential, MLP, and Uniform noise distributions. We find the following interesting outcome: the mixed-mixed architecture is on par with the others on the Sprinkler and the Sachs datasets and outperforms the other methods on the AutoMPG and the Tübingen pairs datasets. Despite these results must be taken cautiously, they provide evidence of a strong result, that mixed training appears to be beneficial even in real-world scenarios, those of actual interest in applications.
Benchmark with classic causal discovery.
We probe CSIvA test generalization in comparison with DirectLiNGAM, CAM, NoGAM, and GraNDAG methods. According to Fig. 7, interestingly we find that the mixed-mixed CSIvA model (trained on SCMs with linear, additive nonlinear and post-nonlinear mechanisms, and Beta, Gamma, Gumbel, Exponential, MLP, Uniform noise distributions) matches with or outperforms the other methods on all the test tasks. This provides additional empirical evidence on the benefits of the mixed training procedure we propose to achieve better test generalization.
Figure 6: Average SHD (the lower, the better) on real-world datasets of CSIvA models that are trained on synthetic datasets generated with linear, nonlinear additive, and post-nonlinear mechanisms and fixed mlp noise distribution (linear-mlp, anm-mlp, pnl-mlp bars) and mixed mechanisms and mixed noise distributions (mixed-mixed bar). Performance is tested on bivariate models. We observe that the model optimized with mixed training is on par or outperforms the other algorithms.
Figure 7: Average SHD (the lower, the better) on real-world datasets. The CSIvA model is trained on synthetic datasets generated with mixed mechanisms and mixed noise distributions (csiva (mixed) bar). As benchmark methods, we consider DirectLiNGAM, CAM, NoGAM, and GraNDAG. Performance is tested on bivariate models. We observe that the model optimized with mixed training is on par or outperforms the other algorithms.
D.2 Benchmarking CSIvA generalization with classical causal discovery algorithms
In this section, we analyze the results of Fig. 8, where we compare the CSIvA trained on mixed mechanisms (linear, nonlinear, post-nonlinear) and mixed noises (all noise except for Gaussian) with the benchmark methods DirectLiNGAM, CAM, NoGAM, GraNDAG. Given that we want to probe CSIvA test generalization, we run inference over the following dataset configurations: linear-mixed (i.e. SCMs considering all possible noise distributions, except for Gaussian), anm-mixed, pnl-mixed, and mixed-mlp (i.e. SCMs with linear, nonlinear, post-nonlinear mechanisms). Fig. 8 shows results in line with our expectations: DirectLiNGAM, CAM, NoGAM, and GraNDAG achieve their best accuracy on data generated by SCMs respecting their assumptions, while degrading their performance on the other models; the CSIvA architecture trained on a mixture of SCMs with different mechanisms and noise distributions matches with or tops all other methods, in all the considered settings (while being outperformed on the anm and linear data, CSIvA still retains good average SHD accuracy).
Figure 8: Average SHD (the lower, the better) on simulated datasets. The CSIvA model is trained on synthetic datasets generated with mixed mechanisms and mixed noise distributions (csiva (mixed) bar). As benchmark methods, we consider DirectLiNGAM, CAM, NoGAM, and GraNDAG. Performance is tested on bivariate models. We observe that, in general, the model optimized with mixed training is on par or outperforms the other algorithms.
D.3 Experiments with different sizes of the training dataset
In this section, we explore how CSIvA test generalization scales when training occurs on different numbers of training samples. In the experiments on the main manuscript, each algorithm is optimized on $15000$ datasets, where each dataset and the underlying causal graph corresponds to a training data point. We now compare the test SHD when training occurs on $5000$ and $10000$ datasets. One clear point emerges from the results of Fig. 9, that is our results on the benefits of the mixed training procedure are consistent for each size of the training dataset we considered. Moreover, we note that the performance of CSIvA does not appear to degrade due to the decrease in the number of training points.
(a) LiNGAM training data (mlp noise)
(b) ANM training data (mlp noise)
(c) PNL training data (mlp noise)
(d) Mixed mechanisms and noise training data
Figure 9: Average SHD (the lower, the better) of CSIvA models trained with $5000$, $10000$, $15000$ data points. The algorithms are tested on simulated datasets generated with linear, nonlinear and post-nonlinear mechanisms (linear-mixed, anm-mixed, pnl-mixed entries on the x axis) and mixed-mlp datasets, generated with mixed mechanism types and fixed mlp noise distribution. We observe that (i) the mixed training improves the test generalization, irrespective of the training dataset size; (ii) CSIvA maintains its performance stable across different training dataset sizes.
D.4 Can we learn to infer causal order from linear Gaussian data?
We ask whether CSIvA trained on non-identifiable models can implicitly learn to predict the causal direction of identifiable SCMs. For this purpose, we consider CSIvA optimized on linear Gaussian data and test its performance on several datasets sampled from structural causal models with different configurations of mechanisms and noise distributions: linear-mixed (with noise terms sampled according to all distributions except for Gaussian), anm-mixed, pnl-mixed, mixed-mlp (with mechanisms generated according to linear, nonlinear, post-nonlinear equations), anm-gauss, and pnl-gauss. The results of Fig. 10 present strong evidence that models trained on non-identifiable SCMs can not infer the causal order: in fact, we see that consistently across all datasets CSIvA average SHD approximates $0.5$, the performance of classification with a coin flip. This is in line with our expectations. In agreement with our Hypothesis 1, in Section 4.3 we have empirically shown that CSIvA can model the class of the SCM generating the observed data and exploit this information to infer the correct causal DAG (instead of a less specific Markov equivalence class) when this is identifiable. Moreover, in our Section 4.2 our experiments show that CSIvA can not generalize to SCM classes unseen during training. In light of these findings, it is intuitive that an architecture trained on non-identifiable linear Gaussian data can only try to fit a linear Gaussian model, irrespective of the input data. Then, when inferring the causal direction, given that CSIvA assumes the data to be generated according to a linear Gaussian SCM, both the forward and backward directions are equally plausible, which explains the observed SHD close to $0.5$.
Figure 10: Average SHD (the lower, the better) of CSIvA trained on datasets generated by linear Gaussian models, which are non-identifiable. Performance is tested on simulated datasets generated according to several SCM configurations. We observe that training on non-identifiable data yields an algorithm that performs with average accuracy of $0.5$, equivalent to a coin flip random baseline, across all the test tasks.
D.5 Experiments with bivariate independent graphs
In the main manuscript, we consider training and testing of CSIvA on bivariate graphs with an edge: $X\rightarrow Y$, $Y\rightarrow X$. This can be phrased as a classification problem with two labels. We motivate our choice by noticing that, in the bivariate setting, identifiability is a property of connected graphs: the empty graph with no edge defines a Markov equivalence class with one element, i.e. a singleton. This is known to be identifiable without explicit assumptions on the functional form of the mechanisms or the noise term distributions in the causal model. The goal of this section is to show that, if we include datasets generated according to an empty graph in the training procedure, CSIvA can learn to disambiguate between the three classes (the empty graph, $X\rightarrow Y$, $Y\rightarrow X$). To motivate our claim, we notice that classifying empty and connected graphs can be done by testing independence between the input variables: previous works phrase independence testing as a classification task and show that this can be learned via deep neural networks (Bellot & van der Schaar, 2019, Sen et al., 2017). The experiments of Fig. 11 sustain our claim. We consider three CSIvA architectures, each trained on independent pairs and one between linear, nonlinear, or post-nonlinear data. Our results show that the neural network can learn to disambiguate between the three classes in all scenarios.
Figure 11: Average SHD (the lower, the better) for CSIvA trained on independent pairs and one between linear, nonlinear, or post-nonlinear data. Each algorithm is tested on the same class of structural causal models it was trained on. We note that in all three scenarios, CSIvA learns to distinguish all classes with almost optimal accuracy (i.e., SHD close to 00.
D.6 Mixed training with unlimited budget
@@
We present our experimental results on one further question, to help clarify the results in the main text of the paper. We aim to understand when to make tradeoffs between computational resources, and having models that have been trained on a wider variety of SCMs. We compare training on multiple SCMs to single-SCM training, when all models see the same amount of training data from each SCM type as a non-mixed model (i.e. a mixed network trains on $15,000$ linear datasets and $15,000$ PNL datasets, instead of $15,000$ divided between the two SCM types).
@@
In the main text of this paper, we compare neural networks trained on a mix of structural causal models (e.g. noise distributions, or mechanism types), to models trained on a single mechanism-noise combination, where all models have the same amount of training data, $15,000$ datasets. In mixed training, we split these evenly, so a "lin, nl" model is trained on $7,500$ datasets from linear SCMs, and $7,500$ from nonlinear SCMs. Our results in this framework are promising, and show that for many combinations of SCM types, we can train one model instead of two, and achieve good progress, while making a $50\%$ savings on training costs. However, if our training budget is high/unlimited, we should also ask whether we can we achieve the same performance as a model trained on a single SCM type. Fig. 12 shows good results in this direction - the models trained with the same number of datasets per SCM type as an unmixed model had similar (or even better, for PNL data) performance as the un-mixed model trained on the same SCM type as the test data. These mixed models are also significantly more useful than having 2 or 3 separate models per SCM type, as they have good across-the-board performance. However, if we used the same computational resources to train 3 separate networks (one for each mechanism type) and wanted to use them for causal discovery on a dataset with unknown assumptions, we would be left with the rather difficult task of deciding which model to trust.
@@
D.7 Experiments with Gaussian process nonlinear mechanisms
In this section we present results obtained training and testing CSIvA on synthetic data with nonlinear mechanisms sampled from a Gaussian process with a unit bandwidth RBF kernel (we call data generated according to this approach as GP-data). In particular, for each variable $X_{i}$ node of the graph $\mathcal{G}$ generated according to model (1) we define the nonlinear mechanism $f_{i}(X_{\operatorname{PA}_{i}^{\mathcal{G}}})=\mathcal{N}(\mathbf{0},K(X_{\operatorname{PA}_{i}^{\mathcal{G}}},X_{\operatorname{PA}_{i}^{\mathcal{G}}}))$, a multivariate normal distribution centered at zero and with covariance matrix as the Gaussian kernel $K(X_{\operatorname{PA}_{i}^{\mathcal{G}}},X_{\operatorname{PA}_{i}^{\mathcal{G}}})$, where $X_{\operatorname{PA}_{i}^{\mathcal{G}}}$ are the observations of the parents of the node $X_{i}$. Together with our strategy adopted in the experiments in the main text of parametrizing nonlinearities with random neural networks (NN-data), this is one of the most common approaches in the literature.
GP-data generation: literature review.
We present an extensive list of works adopting Gaussian processes for the sampling of nonlinear mechanisms: Rolland et al. (2022), Montagna et al. (2023a; b; c), Bühlmann et al. (2014), Mooij et al. (2016), Lachapelle et al. (2020), Wang et al. (2021), Chen et al. (2023), Mooij et al. (2011), Monti et al. (2019). This suggests that our data generation strategy is established in the causality literature.
Summary of the GP-data experiments.
Figures from 13 to 16 replicate the main text experiments involving nonlinear mechanisms either in the training or testing data. The results on GP-data agree with our findings on NN-data: CSIvA still shows poor OOD generalization under different training and test mechanisms, and generally for different training and testing noise distribution (except for PNL data). Similar to the case with NN-data, test generalization improves under mixed training.
(a) Linear
(b) Nonlinear
[…]
Figure 13: In-distribution generalization (GP-data) of CSIvA trained and tested on data generated according to the same structural causal models, fixing mechanisms, and noise distributions between training and testing. Nonlinear mechanisms for nonlinear and pnl data are sampled from a Gaussian process. As baselines for comparison, we use DirectLiNGAM on linear SCMs and NoGAM on nonlinear ANM (we use their causal-learn and dodiscover implementations). CSIvA performance is clearly non-trivial and generalizing well.
(a) Mechanisms
(b) Noise distributions
Figure 14: Out-of-distribution generalisation (GP-data). We train three CSIvA models on data sampled from SCMs with linear, nonlinear additive, and post-nonlinear mechanisms; and fixed mlp noise distribution. Nonlinear mechanisms for nonlinear and pnl data are sampled from a Gaussian process. In Figure 14(a) we test across different mechanism types, with mlp-distributed noise terms both in test and training. In Figure 14(b) we test across different noise distributions, with test mechanism types fixed from training. CSIvA struggles to generalize to unseen causal mechanisms and often displays degraded performance over new noise distributions.
(a) Linear test data
(b) Nonlinear test data
(c) PNL test data
Figure 15: Mixture of causal mechanisms (GP-data). We train four models on samples from structural casual models with different mechanism types. Nonlinear mechanisms for nonlinear and pnl data are sampled from a Gaussian process. We compare their test SHD (the lower, the better) against networks trained on datasets generated according to a single type of mechanism. The dashed line indicates the test SHD of a model trained on samples with the same mechanisms as test SCM. Training on multiple causal models with different mechanisms (mixed bars) always improves performance compared to training on single SCMs.
(a) Linear test data
(b) Nonlinear test data
(c) PNL test data
Figure 16: Mixture of noise distributions (GP-data). We train three networks on samples from SCMs with different noise terms distributions and fixed mechanism types: linear, nonlinear, and post-nonlinear. Nonlinear mechanisms for nonlinear and pnl data are sampled from a Gaussian process. We present their test SHD (the lower, the better) on data from SCMs with the mechanisms fixed with respect to training, and noise terms changing between each dataset. Training on multiple causal models with different noises (all distributions bars) always improves performance compared to training on single SCMs with fixed mlp noise (only mlp bars).
Appendix E Theoretical results and proofs
In this section, we state and prove the identifiability of the post-ANM discussed in Section 4.3, as a corollary of Theorem 1 of Hoyer et al. (2008). The forward and backward models of equations (5) and (6) for the pair of random variables $X,Y$ is given by:
$Y=f_{2}(f_{1}(X)+N_{Y})=f_{2}(\tilde{Y}),\hskip 10.00002pt\tilde{Y}\coloneqq f_{1}(X)+N_{Y},$
$X=g_{2}(g_{1}(Y)+N_{X})=g_{2}(\tilde{X}),\hskip 10.00002pt\tilde{X}\coloneqq g_{1}(Y)+N_{X},$
with $f_{2},g_{2}$ invertible functions, $N_{Y},X$ independent random variables, and $N_{X},Y$ independent random variables. Equivalently, we can frame forward and backward causal models for $\tilde{X},\tilde{Y}$, as in equations (7) and (8):
$\tilde{Y}=h_{Y}(\tilde{X})+N_{Y},\hskip 10.00002pth_{Y}\coloneqq f_{1}\circ g_{2},$
$\tilde{X}=h_{X}(\tilde{Y})+N_{X},\hskip 10.00002pth_{X}\coloneqq g_{1}\circ f_{2}.$
We are now ready to provide our identifiability statement for post-ANMs.
Proposition 1 (Corollary of Theorem 1 of Hoyer et al. (2008)).